OpenStudy (anonymous):

By using 70.0 J of work, a black of wood is moved 2.4m across that was applied to the block. Honestly I have no Idea what this is talking about or how to do it.

OpenStudy (anonymous):

could you please retype task...

OpenStudy (anonymous):

what do you need to calculate?! Force or what?

OpenStudy (anonymous):

I do believe force but I have no Idea truly, I was given this question on my review for the final exams and I don't have a clue what it's asking

OpenStudy (anonymous):

W = F * s W - work F - force s - distance W - J - Joule - Nm F - N - Newton s - m -meter F= W/s = 70 J / 2,4 m = 29.16 N